# ORFS-agent: Tool-Using Agents for Chip Design Optimization

### v1

@@ Appendix B Additional Experimental Details
B-A Details on Parameters, Metrics Tracked, and Surrogates
The details of parameters are shown in Table VII. Table VIII summarizes metrics tracked and surrogates.
@@ B-A Details on Parameters, Metrics Tracked, and Surrogates
TABLE VII: Overview of optimization parameters for both 4- and 12-parameter tuning. The top four rows in bold form the 4-parameter case.
@@
DRWL ($WL$) Post-DR wirelength.
@@
EPC (PDP) Energy per clock cycle, a.k.a. Power Delay Product (PDP)
@@
The Clock Period is specified to one decimal place. Core Utilization and TNS End Percent are integers between 00 and $100$ inclusive, while LB Add On Place Density (henceforth LB Add On) is a real number specified to two decimal places between 00 and $1$. Thus, the latter three variables have $100$ discrete values each, while the Clock Period remains unbounded.
The Effective Clock Period is defined as the clock period minus the worst slack at the finish stage. These metrics, along with their CTS-level surrogates and other desirable figures of merit, are defined in Table VIII. These are based on the METRICS2.1 suite. The Optimized metrics directly enter the natural language optimization process for the LLM. The Surrogates are included only if the primary metrics are unavailable due to timeouts. The remaining parameters—Instance Area, Instance Count, PDP, Power—are not directly optimized but are tracked as meaningful quantities. In the expanded separate configuration (“12-variable setting”), we introduce eight additional tunable parameters that guide the physical design flow more comprehensively. Specifically:
@@
Appendix C Analysis and Ablations under ASAP7 IBEX
Effects of Iterations and Indirect Optimization. We will consider the results of the joint $WL,ECP$ optimization as a function of the total serial iterations (keeping the number of parallel runs fixed at $25$). The results are shown in Table IX, along with the other relevant figures of merit from Table VIII. Note that each row represents a separate run that is run upto the corresponding number of serial iterations with $25$ parallel runs. That is, $5$ should be considered as a total of $25\times 5=125$ OpenROAD calls.
@@ Appendix C Analysis and Ablations under ASAP7 IBEX
TABLE IX: Optimization trajectory $5-30$ iterations, no tool use, 4 parameters.
This phenomenon can be well understood in the light of the strong inter-correlations of the data, presented in the main text.
Reversibility. From Table IX, it is clear that optimizing wirelength and Effective Clock Period (ECP) jointly also optimizes other parameters. We now reverse this procedure to analyze the effect on wirelength and ECP when any one of these parameters is optimized individually. The results of this process are presented in Table X. Notably, this process appears to be asymmetric—these iterations yield noticeably worse outcomes on ECP while resulting in significantly lower power consumption.
@@
TABLE X: Results of optimizing the row-level objective on column-level outcomes, ASAP7 IBEX, no tool use, 4 parameter setting.
To ensure that our results stay consistent across tool use choices, we replicate this experiment on ASAP7 IBEX with 12 parameters and tool use enabled in Table XI.
@@
TABLE XI: Results of optimizing the row-level objective on column-level outcomes, ASAP7 IBEX, tool use and 12 parameter setting.
Constrained Optimization under Natural Language Commands. We presented the results of the constrained optimization setup in the main text. Here, we continue on this vein. A human engineer is typically given commands significantly different from the utility functions used by OR-AutoTuner. Specifically, a common class of optimization commands relayed to humans takes the form:
“Optimize metric $X$ as much as possible, while not allowing metrics $A,B,C$ to deteriorate by more than $2\%$.”
This form of constrained optimization is generally not straightforward with existing OR-AutoTuner setups. LLMs can attempt such optimization by incorporating these constraints directly into their context. We study this phenomenon in Table XII and compare it to OR-AutoTuner in Tables XIII and  XIV, where we optimize $ECP$ and $WL$ subject to a maximum worsening of $2\%$ and $4\%$ in other metrics of interest—Area, Instance Count, Power, and PDP. These percentage values are always with respect to the baseline values generated from default OR-AutoTuner parameters.
@@
TABLE XII: Constrained optimization results. Red positive numbers indicate worsening, i.e. an increase relative to baseline, and the “leeway” indicates the maximum worsening allowed.
@@
TABLE XIII: Constrained optimization results for OR-AutoTuner (with 375 iterations, four parameters). Red positive numbers indicate worsening, i.e., an increase relative to baseline, and the “leeway” indicates the maximum worsening allowed.
@@
TABLE XIV: Constrained optimization results for OR-AutoTuner (with 1000 iterations, 12 parameters). Red positive numbers indicate worsening, i.e., an increase relative to baseline, and the “leeway” indicates the maximum worsening allowed.
The $4\%$ tolerance case naturally gives rise to a Pareto frontier scatterplot of Effective Clock Period (ECP) versus wirelength (WL), which is shown in Figure 5.
Figure 5: Pareto frontier of ECP vs. WL under $4\%$ tolerance, relative to the baseline. The blue line represents the case with 4 parameters and no tool use, and the red line represents the case with 12 parameters and tool use.
@@ Appendix D Ablations for Data Leakage, Confidence and Robustness
We begin by analyzing statistical significance in table XV, and testing prompt-level changes in tables XVI and  XVII. The “role” here should be understood as the role of an EDA expert that the LLM is asked to pose as. The details exist separately on the level of PDK / platform, circuit, and parameters. Each of these is ablated against. The reader should carefully also check the implementation codebase provided with the paper to understand the nature of this metadata.
TABLE XV: Statistical significance over 20 independent ORFS-agent runs on the ASAP7 IBEX circuit. Percentile columns report the $k$-th worst value ($k\!=\!1,5,10,15,19$) among the 20 outcomes; “Reported” is the single-run value quoted in the main paper.
Metric $5$-th $25$-th Median $75$-th $95$-th Reported
@@
TABLE XVI: Ablation on domain-knowledge content in prompts for ASAP7 IBEX. “PDK” removes technology-node details; “circuit” removes design-specific details; “Both” drops both. Higher WL/ECP indicate worse quality of result.
@@
TABLE XVII: Sensitivity of ORFS-agent to prompt simplifications on ASAP7 IBEX. “Removed role” deletes the system-role description; “Removed param” deletes parameter-guidance text.
@@
Throughout the process, the large amount of data from Anthropic raises concerns about whether the LLM is simply caching or memorizing what it observes. This includes the possibility of interactions with other OpenROAD users who may be testing Claude models on similar outputs. We test these concerns in the following ways:
Circuit Obfuscation. To assess potential memorization, we obfuscate the name of the ASAP7-IBEX circuit and subject it to the same iteration vs. performance ablation test. The circuit name is changed to obfs, while the design platform input to the ORFS-agent remains ASAP7. Additionally, configuration settings are modified to rename the top-level module from “IBEX” to obfs. The ORFS-agent is also restricted from grepping log outputs that refer to the original name or netlist identifiers. The results appear in Table XVIII and Table XIX.
@@
TABLE XVIII: Obfuscated results—ASAP7 IBEX, no tools and with 4 parameters.
@@
TABLE XIX: Obfuscated results—ASAP7 IBEX, with tools and with 12 parameters, including PDP = ECP $\times$ Power.
Statistical Significance. We run the ASAP7-IBEX optimization for $ECP,WL$ $10$ times, with results in Table XX and  XXI. Even the worst of the $10$ produces a result that is superior in terms of ECP against OR-AutoTuner. Each individual row indicates the outcome with the lowest ECP in the final $25$ iterations of an individual run.
@@
TABLE XX: $10$-fold runs, ASAP7 IBEX, 4 parameters and no tools.
@@
TABLE XXI: 10-fold runs, ASAP7 IBEX, 12 parameters and with tools.
Robustness Testing. We perform two layers of robustness testing. First, we add noise to every measurement: we multiply every observation $O$, at random, with the variable $1+X_{O}$, where each $X_{O}$ is an i.i.d. uniform random variable from $[-0.05,0.05]$. Next, we consider the case where our specification of the surrogate has gone astray - it is a common occurrence in EDA for the variable being optimized to not really track what is desired. For this, we change all the CTS-level observations to be multiplied by $1+X_{CTS}$, where $X_{CTS}$ is an uniform random variable i.i.d drawn from $[-0.2,0.2]$. The results on ASAP7 IBEX are presented below in Table XXII.
@@
WL & ECP 109873 1281 2705 21877 93.51 0.073
TABLE XXII: Perturbed runs, 12 parameters, tool use permitted.
@@ Appendix E Global Context, Initialization, and the Toolbox
LLM calls made within each part of an iteration - i.e., INSPECT, etc. - add the LOCALCONTEXT of that particular iteration. However, they are not necessarily passed on to the GLOBALCONTEXT. As we move from iteration $j$ to $j+1$, GLOBALCONTEXT is modified to only include a small subset of all the LOCALCONTEXT within the particular iteration $j$.
@@ E-A Initialization
  • The outputs, specifying the variable sets (e.g., detailed WL) and variable sets of its surrogate(s) (e.g., CTS WL).
@@
  • The domains of the inputs, including whether they are Boolean, integers with min/max bounds, or continuous values converted to integers (e.g., converting float ranges like 0.01–0.99 to integer ranges 1–99, and scaling back to original values for PARAMETERS generation).
@@ E-B Toolbox
We define three separate kinds of tools. All of them require a Description, a set of ideally typed arguments, and a return object Return with description and type specifications to be efficiently made a part of LLM function-calling frameworks. All of them are summarized in section F.
@@
OPTIMIZE Tools. These tools propose new points (parameter sets) to explore in an optimization process. Each one uses a different strategy—ranging from random sampling and grid-based approaches to more sophisticated algorithms like Bayesian optimization or genetic algorithms—to balance exploration and exploitation.
@@ Appendix F Tool Details
CreateModel creates and configures a Gaussian Process model. It takes X: np.ndarray, y: np.ndarray, noise_level: float, and kernel_type: str, returning a model of type GaussianProcessRegressor.
CreateKernel generates a kernel from a given specification. The required inputs are kernel_spec: str and input_dim: int. It outputs a GPR kernel.
ExpectedImprovement calculates the Expected Improvement (EI) acquisition function. Inputs include mu: np.ndarray, std: np.ndarray, and y_best: float. The return is ei: np.ndarray.
HandleSurrogate processes and combines true and surrogate data. It uses X: np.ndarray, y: np.ndarray, and surrogate_values: np.ndarray, and returns a tuple (X, y_combined, uncertainty).
LatinHypercube generates Latin Hypercube samples. The inputs are n_points: int and n_dims: int, returning samples: np.ndarray.
@@
SelectPoints selects points based on a specified method. It accepts X: np.ndarray, quality_scores: np.ndarray, method: str, and n_points: int, and returns selected_indices: np.ndarray.
HybridSelect combines quality and diversity for point selection. Required arguments are X: np.ndarray, quality_scores: np.ndarray, distance_matrix: np.ndarray, and n_points: int, with return selected_indices: np.ndarray.
EntropySelect uses entropy-based diversity for selection. It takes X: np.ndarray, quality_scores: np.ndarray, and n_points: int, returning selected_indices: np.ndarray.
GraphSelect applies graph-based diversity metrics for point selection. It accepts X: np.ndarray, quality_scores: np.ndarray, and n_points: int, returning selected_indices: np.ndarray.
@@
Each of the tools used here can be implemented using standard python packages. For instance, scikit-learn, numpy and related Python packages offer support for Gaussian Processes, t-SNE modeling, and so on. By combining these existing libraries with domain-specific logic, we create a modular, maintainable toolbox that reflects the iterative exploration process a human data scientist would perform —streamlining tasks like outlier detection, dimensionality reduction, and optimization-driven design.

### v2

@@ Appendix B Additional Experimental Details
B-A Parameters, Metrics Tracked, and Surrogates
Table VII gives details of optimization parameters. The Clock Period is specified to one decimal place, and has no upper bound. Core Utilization and TNS End Percent are integers between 0 and 100100100 inclusive, while LB Add On Place Density (henceforth LB Add On) is a real number between 0 and 111, specified to two decimal places.
@@ B-A Parameters, Metrics Tracked, and Surrogates
TABLE VII: Overview of optimization parameters for both 4- and 12-parameter tuning. The top four rows form the 4-parameter case.
@@
$WL$ Post-detailed-routing wirelength.
@@
EPC (PDP) Energy per clock cycle equal to Power–Delay Product (PDP)
@@
Table VIII summarizes metrics tracked along with surrogates. The Effective Clock Period is defined as the Clock Period minus the worst slack at the finish stage. The metrics in Table VIII, along with their CTS-level surrogates and other desirable figures of merit, are based on the METRICS2.1 suite. The Optimized metrics directly enter the natural language optimization process for the LLM. The Surrogates are included only if the primary metrics are unavailable due to timeouts. The remaining parameters – Instance Area, Instance Count, PDP, Power – are not directly optimized but are tracked as meaningful quantities. The expanded “12-variable setting” configuration adds the following eight tunable parameters that more comprehensively guide the physical design flow.
@@
Appendix C Analyses and Ablations with ASAP7-IBEX
Effects of iterations and indirect optimization. We consider the results of the joint $WL,ECP$ optimization as a function of the total serial iterations (keeping the number of parallel runs fixed at 252525). The results are shown in Table IX, along with the other relevant figures of merit from Table VIII. Note that each row represents a separate run that is executed up to the corresponding number of serial iterations with 252525 parallel runs. E.g., 555 should be considered as a total of 25×5=12525$\times$ 5=12525 × 5 = 125 OpenROAD calls.
@@ Appendix C Analyses and Ablations with ASAP7-IBEX
TABLE IX: Optimization trajectory with 5 to 30 iterations, no tool use, 4-parameter setting.
This phenomenon can be well understood in the light of the strong inter-correlations of the data, presented in the main text (Subsection IV-B).
Reversibility. Table IX, indicates that optimizing wirelength and $ECP$ jointly also optimizes other parameters. We also reverse this, and analyze the effect on wirelength and $ECP$ when any one of these other parameters is optimized individually. The results of this process are presented in Table X. Notably, outcomes are asymmetric in that we see noticeably worse $ECP$ outcomes, but with significantly lower power consumption.
@@
TABLE X: Results of optimizing the row-level objective on column-level outcomes: ASAP7-IBEX, no tool use, and 4-parameter setting.
To confirm that our results are consistent across tool use choices, we replicate this experiment using ASAP7-IBEX with 12 parameters and tool use enabled; results are presented in Table XI.
@@
TABLE XI: Results of optimizing the row-level objective on column-level outcomes: ASAP7-IBEX, tool use, and 12-parameter setting.
Constrained optimization under natural language commands. We have presented the results of the constrained optimization setup in Subsection IV-D of the main text. Here, we continue on this vein. A human engineer is typically given commands significantly different from the utility functions used by OR-AutoTuner. Specifically, a common class of optimization commands relayed to humans takes the form:
[…]
@@
TABLE XII: Constrained optimization results. Red positive numbers indicate worsening, i.e., an increase relative to baseline; “leeway” indicates the maximum worsening allowed.
@@
TABLE XIII: Constrained optimization results for OR-AutoTuner with 375 iterations, four parameters. Red positive numbers indicate worsening, i.e., an increase relative to baseline; “leeway” indicates the maximum worsening allowed.
@@
TABLE XIV: Constrained optimization results for OR-AutoTuner with 1000 iterations, 12 parameters. Red positive numbers indicate worsening, i.e., an increase relative to baseline; “leeway” indicates the maximum worsening allowed.
[…]
@@ Appendix D Ablations for Data Leakage, Confidence and Robustness
In this appendix, we analyze statistical significance in Table XV, and study prompt-level ablations in Tables XVI and  XVII. Table XVI presents outcomes of separate ablations of PDK (i.e., platform), circuit, and parameters. We refer readers to the implementation codebase [80] to understand the nature of this metadata. It can be clearly seen that the median values reflect the conclusions in Section IV and Section V of the main body of the paper, and that the interquartile ranges are small enough relative to the median to allow some assurances as to robustness.
Table XVII shows additional ablations, where “role” refers to the role of an EDA expert that the LLM is asked to pose as.
Metric $5^{th}$ $25^{th}$ Median $75^{th}$ $95^{th}$ Reported
@@
TABLE XV: Statistical significance over 20 independent ORFS-agent runs on the ASAP7-IBEX circuit. Percentile columns report the $k^{th}$ worst value ($k\!=\!1,5,10,15,19$) among the 20 outcomes; “Reported” is the single-run value quoted in Section IV of the main paper.
@@
TABLE XVI: Ablation on domain-knowledge content in prompts for ASAP7-IBEX. “No PDK” removes technology-node details; “No circuit” removes design-specific details; “Both removed” drops both. Higher $WL$/$ECP$ indicate worse quality of result.
@@
TABLE XVII: Sensitivity of ORFS-agent to prompt simplifications on ASAP7-IBEX. “No role” deletes the system-role description; “Blank param” deletes parameter-guidance text.
Over the course of our work, the large amount of data from Anthropic naturally raises concerns about whether the LLM is simply caching or memorizing what it observes. This includes the possibility of interactions with other OpenROAD users who may be testing Claude models on similar outputs. We test these concerns in the following ways.
Circuit obfuscation. To assess potential memorization, we obfuscate the name of the ASAP7-IBEX circuit and subject it to the same iteration vs. performance ablation test. The circuit name is changed to obfs, while the design platform input to the ORFS-agent remains ASAP7. Additionally, configuration settings are modified to rename the top-level module from “IBEX” to obfs. The ORFS-agent is also restricted from grepping log outputs that refer to the original name or netlist identifiers. Results are given in Tables XVIII and  XIX.
@@
TABLE XVIII: Obfuscated results for ASAP7-IBEX, no tools, 4 parameters.
@@
TABLE XIX: Obfuscated results for ASAP7-IBEX, with tools, 12 parameters.
Statistical significance. We run the ASAP7-IBEX optimization for $ECP$, $WL$ 101010 times, with results given in Tables XX and  XXI. Each individual row indicates the outcome with the lowest $ECP$ in the final 252525 iterations of an individual run. Even the worst of the 101010 produces a result that is superior to OR-AutoTuner in terms of $ECP$.
@@
TABLE XX: 101010-fold runs, ASAP7-IBEX, no tools, 4 parameters.
@@
TABLE XXI: 10-fold runs, ASAP7-IBEX, with tools, 12 parameters.
Robustness testing. We perform two layers of robustness testing. First, we add noise to every measurement: we multiply every observation $O$, at random, with the variable $1+X_{O}$, where each $X_{O}$ is an i.i.d. uniform random variable drawn from [0.05,0.05][-0.05,0.05][ - 0.05 , 0.05 ]. Next, we consider the scenario where our specification of the surrogate has gone astray (i.e., the objective being optimized does not track what is desired. For this, we change all the CTS-level observations to be multiplied by $1+X_{CTS}$, where $X_{CTS}$ is an i.i.d. uniform random variable drawn from [0.2,0.2][-0.2,0.2][ - 0.2 , 0.2 ]. The results for ASAP7-IBEX are presented in Table XXII.
@@
$WL$ $ECP$ 109873 1281 2705 21877 93.51 0.073
TABLE XXII: Perturbed runs, with tools, 12 parameters.
@@ Appendix E Global Context, Initialization, and the Toolbox
LLM calls made within each part of an iteration – i.e., INSPECT, etc. – add the LOCALCONTEXT of that particular iteration. However, they are not necessarily passed on to the GLOBALCONTEXT. As we move from iteration $j$ to iteration $j+1$, GLOBALCONTEXT is modified to only include a small subset of all the LOCALCONTEXT within the particular iteration $j$.
@@ E-A Initialization
  • The outputs, specifying the variable sets (e.g., detailed $WL$) and variable sets of surrogate(s) (e.g., CTS $WL$).
@@
  • The domains of the inputs, including whether they are Boolean, integers with min/max bounds, or continuous values converted to integers (e.g., converting float ranges such as 0.01–0.99 to integer ranges such as 1–99, and scaling back to original values for PARAMETERS generation).
@@ E-B Toolbox
We define three separate kinds of tools. All of these require a Description, a set of ideally typed arguments, and a return object Return with description and type specifications to be efficiently made a part of LLM function-calling frameworks. We summarize tool details in Appendix F.
@@
OPTIMIZE tools. These tools propose new points (parameter sets) to explore in an optimization process. Each one uses a different strategy – ranging from random sampling and grid-based approaches to more sophisticated algorithms such as Bayesian optimization or genetic algorithms – to balance exploration and exploitation.
@@ Appendix F Tool Details
CreateModel creates and configures a Gaussian process model. It takes X: np.ndarray, y: np.ndarray, noise_level: float, and kernel_type: str. It returns a model of type GaussianProcessRegressor.
CreateKernel generates a kernel from a given specification. Required inputs are kernel_spec: str and input_dim: int. The output is a GPR kernel.
ExpectedImprovement calculates the Expected Improvement (EI) acquisition function. Inputs include mu: np.ndarray, std: np.ndarray, and y_best: float. It returns ei: np.ndarray.
HandleSurrogate processes and combines true and surrogate data. It uses X: np.ndarray, y: np.ndarray, and surrogate_values: np.ndarray. It returns a tuple (X, y_combined, uncertainty).
LatinHypercube generates Latin Hypercube samples. The inputs are n_points: int and n_dims: int. It returns samples: np.ndarray.
@@
SelectPoints selects points based on a specified method. It accepts X: np.ndarray, quality_scores: np.ndarray, method: str, and n_points: int. It returns selected_indices: np.ndarray.
HybridSelect combines quality and diversity for point selection. Required arguments are X: np.ndarray, quality_scores: np.ndarray, distance_matrix: np.ndarray, and n_points: int. It returns selected_indices: np.ndarray.
EntropySelect uses entropy-based diversity for selection. It takes X: np.ndarray, quality_scores: np.ndarray, and n_points: int. It returns selected_indices: np.ndarray.
GraphSelect applies graph-based diversity metrics for point selection. It accepts X: np.ndarray, quality_scores: np.ndarray, and n_points: int. It returns selected_indices: np.ndarray.
@@
Each of the tools used here can be implemented using standard Python packages. For instance, scikit-learn, numpy and related Python packages offer support for Gaussian processes, t-SNE modeling, and so on. By combining these existing libraries with domain-specific logic, we create a modular, maintainable toolbox that reflects the iterative exploration process a human data scientist would perform – streamlining tasks such as outlier detection, dimensionality reduction, and optimization-driven design.
Appendix G Results with better LLMs
Recently, Anthropic released the Claude 4[67] class of models – Sonnet and Opus 4. Our main results use Sonnet 3.5, and hence our ongoing work seeks to rigorously quantify the benefit from more recent models.
[…]
  • Clock Period: 5764 ps
  • Core Utilization: 69.0 %
  • TNS End Percent: 100
  • Global Padding: 1
  • Detail Padding: 0
  • Enable DPO: 1
  • Pin Layer Adjust: 0.48
  • Above Layer Adjust: 0.10
  • Density Margin Add-On: 0.384
  • Flatten Hierarchy (HIER_SYNTH): 1
  • CTS Cluster Size: 28.0
  • CTS Cluster Diameter: 80.0
The worst slack in this case is 161216121612ps, resulting in an $ECP$ of 415241524152ps, which is worse than the baseline $ECP$. This may indicate an ability of Sonnet 4 models to find solutions that degrade QoRs outside the optimization target, which we did not observe in Sonnet 3.5 models.